OpenStudy (anonymous):

Pre Algebra Help!! HELP ASAP!! 1)4^5/4^5 2)120m^16/45m^8 3)8b^3/20b^3 4)3c^8/12c^12 5)-7x^3*8x^2 6)Explain the process you would use to simplify this. 5^3*5^2/5^7 Please help me anyone!!!

OpenStudy (anonymous):

for the first one do u know how to do it?

OpenStudy (anonymous):

ummm well...not really do i just do 4*4*4*4*4/4*4=1024? Correction - the 2nd 4^5 was meant to me 4^2

OpenStudy (anonymous):

This seems to be indices/powers, have you learnt anything called index laws?

OpenStudy (anonymous):

no. i haven't.

OpenStudy (anonymous):

\[a ^{m} \div a ^{n} = a ^{m-n}\] \[a ^{m}\times a ^{n} = a ^{m+n}\]

OpenStudy (anonymous):

those are the basic ones, can you try do them now?

OpenStudy (anonymous):

do they say aw^2*aw^=aw^2+w2 and aw^2*aw^=aw^2-w2??

OpenStudy (anonymous):

can someone please just answer the questions i have listened at the top. thanks.

OpenStudy (anonymous):

1)4^5/4^5 The law of Quotient of Powers says that,\[a^7\div a^4 = a ^{7 - 4}\]When dividing like bases, SUBTRACT the exponents.

OpenStudy (anonymous):

you see, our purpose is to not provide answers for you put rather try to guide and help you.

OpenStudy (anonymous):

@jazy has the right solution for the first one and when it is times remember to add the exponents, refer back to my type basic examples for reference.